OpenStudy (anonymous):

limit. (exponent and lhospital) limit (1-1/x)^(0.7x) as x->infinity (formatting in reply)

OpenStudy (anonymous):

eyeball it

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(1-\frac{ 1 }{ x })^{0.7x}\] Do I take a natural log? I'm not sure how to work this.

OpenStudy (anonymous):

oh, ok, then (1 - 0)^0 ?

OpenStudy (ksaimouli):

power rule

OpenStudy (anonymous):

let me say it correctly \[\lim_{x\to \infty}(1+\frac{1}{x})^x=e\] that is pretty much by definition it follows that \[\lim_{x\to \infty}(1-\frac{1}{x})^x=e^{-1}=\frac{1}{e}\]

OpenStudy (anonymous):

and that \[\lim_{x\to \infty}(1-\frac{1}{x})^{-7x}=e^{-.7}=\frac{1}{e^{.7}}\]

OpenStudy (anonymous):

you can of course take the log, and use l'hopital for that, then exponentiate that will probably make your math teacher happier

OpenStudy (anonymous):

take the log, get \[.7x\ln(1-\frac{1}{x})\] now you are in the form \(\infty\times 0\) usual trick is to rewrite as \[\frac{\ln(1-\frac{1}{x})}{\frac{1}{.7x}}\] then take the derivative top and bottom after a bunch of boring algebra you get \(-.7\) so your answer will be \(e^{-.7}\)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty}(1+\frac{ 1 }{ x })^x = e\] is a known rule that I should remember?

OpenStudy (anonymous):

ok, that last part with the ln and LH looks more familiar, I'm going to work that out for a min.

OpenStudy (anonymous):

what is your definition of \(e\)? sometimes that is used to define the number

OpenStudy (anonymous):

yes, you can take the derivative top and bottom. it will be some annoying algebra maybe easier if you write \[\ln(1-\frac{1}{x})\] as \[\ln(\frac{1-x}{x})\]

OpenStudy (anonymous):

just that e = 2.72 something. wow, this will be a bunch of harsh algebra. How does e get introduced this way? (the ln / algebra route)